Commissioning: 19 May 2010 UT

Observer: Andrew, Dave T
Telescope Operator: SAllanson, WWack (in Tucson)
Telescope Support: Brian (LBTO)
Instrument Support: TShih (LBTO)
Adaptive Secondary Operator: Runa (LBTO)

Summary

Frustrating night, a bit of everything, little progress.

Preparations

7:20 Laser tracker set on c-ring extension; Runa ready to flatten DX M2, our program begins momentarily, outside temperature 2.1 degrees, SX M1 temperature 3.1

7:57 checked with laser tracker and found we had wrong collimation model loaded. proceed to take "binocular test" before going back to that.

7:58 guiding and collimating on SX LFBG BS9166 GS0 87 EL

8:00 note that the GCS GUI reported RMS WFE is inconsistent with listed Zernike terms

8:15 DX M2 was in simulation mode (!!!!!!!), editing conf file and restarting OSS. 8:16 TCS crashed. Calling Norm 8:35 Turns out that a typo "FLASE" instead of "FALSE" in the lbt.conf file is what caused the crash 9:10 With Norm's assistance back online now starting again at the beginning. Runa preparing DX M2.

9:20 Lucifer problem kills this activity (simultaneous imaging on DX and SX FBG), see note on issue track 2463.

9:30, switched off cooling and "stopped" AGW2. Reconfiguring telescope for LBC.

9:40 authorizing as LBC, clearing M2 offloads...

Details

10:00 begin the nights activities at 10:00 UT

10:14 completed laser tracker verification of M1 collimation model. Looks ok at first glance.

10:15 begin collecting pointing model data. Begin by running DOFPIA on high elevation star (85 degrees ACT0302)

11:10 Had a lot of trouble with DOFPIA couldn't get a hole in blue no matter how much Z11 was added. Ended up loading old collimation model and trying again. Converged in 1.7 " seeing. Cleared active optics and Z total -1.817 after instrument offset removed. Reloaded new collimation model and adjusted golbal Z to -1.589.

11:25 took image at 78 degrees elevation, confirmed useful image. (Have given up on tasking a set of pointing data tonight due to lack of time, so are just checking out our procedure for tomorrow night.) LBCB.20100519.112628

11:30 going to ACT 0200 at ~55 degrees EL. Image LBCB.20100519.113141 good images.

11:32 going to ACT 0111 at ~ 27 degrees EL. Image LBCB.20100519.113508 good images.

11:35 going to ACT 0107 at ~ 3 degrees el, Image *.113810, not good but still useable!

11:40 Closing for twilight.

-- DavidThompson - 19 May 2010
Topic revision: r5 - 19 May 2010, JoarBrynnel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback